# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v2

@@ Abstract
Diffusion models have shown remarkable capabilities in modeling complex data distributions by transforming noise into structured data through stochastic processes. However, when applied to dense prediction tasks whose goal is to capture per-pixel relationships between RGB images and dense signal maps, starting the sampling process from an uninformative Gaussian noise often leads to inefficient sampling and long latency. To overcome these challenges, we propose DPBridge, a generative framework that establishes direct mapping between input RGB images and dense signal maps based on a tractable bridge process. Furthermore, we introduce finetuning strategies to leverage a pretrained large-scale image diffusion backbone, enjoying its rich visual prior knowledge to enable both efficient training and robust generalization. Experiments show that DPBridge achieves competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, validating its effectiveness and adaptability.
Machine Learning, ICML
Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing conditional probability transport between RGB images and corresponding dense maps in the latent space based on a tractable diffusion bridge process. In addition, we introduce finetuning strategies to adapt large pretrained diffusion backbone to construct our bridge, leveraging its rich prior knowledge to enable both efficient training and robust generalization.
@@ 1 Introduction
Dense prediction refers to the type of problems whose primary goal is to assign a value at every spatial location in the input data, generating output signal maps that possess same structure and size as inputs. It incorporates various typical tasks, including depth estimation (Yang et al., 2024; Ranftl et al., 2021; Eftekhar et al., 2021), surface normal prediction (Bae & Davison, 2024; Hu et al., 2024; Bae et al., 2021), semantic segmentation (Chen, 2017; Cheng et al., 2022; Wang et al., 2023; Weber et al., 2021), optical flow (Teed & Deng, 2020; Huang et al., 2022; Weinzaepfel et al., 2023), etc.
Recently, image diffusion models have been applied to dense prediction tasks by formulating them as conditional generation problems. Their ability to capture complex data distributions makes them effective priors, leading to highly favorable results (Ke et al., 2024; Garcia et al., 2024; Fu et al., 2025). Following a common practice, these methods diffuse signal maps into pure noise and subsequently leverage input RGB images as conditions to guide the generation process. Nonetheless, there still exists limitations in diffusion-based solutions, mostly stem from the underlying modeling mechanism. The denoising process of a standard diffusion model always starts from Gaussian noise, which contains little information about the target signal distribution. In such cases, the distance between prior Gaussian distribution and target distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms, at the risk of degraded performance and prolonged convergence time.
Instead of starting from noise, one feasible alternative is to directly initiate the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior. This leads to the exploration of a new family of computational frameworks, i.e. diffusion bridge models (Liu et al., 2022; Zhou et al., 2023; Li et al., 2023; Liu et al., 2023). They are based on stochastic process with fixed starting and ending points and are adept at modeling conditional probability paths. By establishing direct connections between different distributions, this line of framework provides a more natural and efficient way to model probability transport compared with previous efforts by reducing the dependence on manually designed post-hoc guidance (Chung et al., 2022a; Song et al., 2023; Chung et al., 2022b) or customized generation pipeline (Meng et al., 2021). We therefore advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
Motivated by this, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model to handle general dense prediction tasks through generative pipeline. In essence, we modify the standard diffusion process to directly bridge the latents of RGB images and corresponding dense signal maps, and enforce the model to learn this process through maximum likelihood estimation. In this way the generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from RGB images rather than pure noise, enabling faster inference speed as well as better generation quality.
Furthermore, we introduce novel finetuning strategies to better adapt pretrained diffusion backbone to DPBridge construction. The equipped rich visual prior not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence. Figure 1 is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e. depth estimation, surface normal prediction, and semantic segmentation, proving that our framework is theoretically universal and can be instantiated to different scenarios.
To summarize, our contributions are as follows:
@@
  • We propose the DPBridge framework to address dense prediction tasks within a generative pipeline by formulating them as instances of latent diffusion bridges.
  • We propose finetuning strategies to effectively adapt large pretrained backbone to diffusion bridges, allowing for exploitation of visual prior knowledge and training efficiency improvements.
@@
  • Our proposed method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection, highlighting its effectiveness and generalization capability.
@@ 2.1 Dense Prediction
Dense predictions can be characterized as a conditional distribution modeling task and common solutions can be categorized into discriminative and generative ones. Discriminative methods directly map input data to structured outputs in a feed-forward manner. (Ranftl et al., 2021) introduces attention mechanism and transformer architecture to better capture both local and global context feature. (Yang et al., 2024) explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness. (Bae & Davison, 2024) incorporates per-pixel ray directions and models the relative rotations between neighboring pixels. Recently, diffusion-based generative models have been applied in this field and showed remarkable performance. (Ke et al., 2024) finetunes pretrained diffusion model, leveraging the rich visual knowledge within to enhance depth estimation. (Fu et al., 2025) takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions. (Gui et al., 2024) exploits flow matching to facilitate the distribution shift between images and depth maps. Our work follow conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in iterative denoising manner.
2.2 Diffusion and Bridge Models
@@
Diffusion models (Ho et al., 2020; Song et al., 2020b) are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes. They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis (Rombach et al., 2022; Ho et al., 2022), controlled editing (Zhang et al., 2023; Batzolis et al., 2021), inverse problem solving (Chung et al., 2022a; Song et al., 2023), 3D assets generation (Xiang et al., 2024; Zhang et al., 2024), and human-centered applications (Ji et al., 2024b, a; Ji & Li, 2024).
@@
Diffusion bridge models have gained increasing attention in generative modeling in recent years (Liu et al., 2022; Zhou et al., 2023; Liu & Wu, 2023; De Bortoli et al., 2021). They offer a novel perspective for translating between different distributions and enable a more effective data-to-data generation paradigm. Recent works have explored different types of bridge processes and their applications. (Liu et al., 2023; Li et al., 2023; Yue et al., 2023) develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution. (Chung et al., 2024) proposes a modified inference procedure that enforces data consistency. (Chen et al., 2023) introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency. To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness. Our work aims to address this research gap.
@@ 3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that is conditioned to end at a specified terminal point at timestep $T$. One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints. We can use Doob’s h-transform technique (Oksendal, 2013; Särkkä & Solin, 2019) to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value. The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
@@
$m_{t}$ $=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right)$
$n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}}$
$\sigma_{t}$ $=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}},$
@@ 4.1 Problem Formulation
We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g. depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image. In the following sections, we’ll elaborate the details of DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
Figure 2: Overview of our training pipeline. To overcome the discrepancy with standard diffusion process and facilitate finetuning, we derive the maximum likelihood training objective of diffusion bridge process, propose a distribution aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
@@ 4.2 Dense Prediction Bridge
Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Section 3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more, it lacks the ability to harness rich visual prior knowledge inherent in the foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks. Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between bridge process and standard diffusion process, aiming for more efficient knowledge transfer. Figure 2 gives an overview of the proposed finetuning pipeline, and next we’ll explain the design of each component.
@@ 4.3 Latent Space Implementation
Instead of directly learning in image pixel space, we follow common practice to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the dense signal maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the initial and terminal points. In this way the diffusion bridge process is henceforth learned in the latent space. To retrieve original signal map we can use the decoder to decode the predicted latents $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. In later part of this paper, with little abuse of notation, all derivations are conducted in the latent space.
@@ 4.4 Training Objective
4.4.1 Maximum Likelihood Loss
Widely-used techniques for learning the reverse process that corresponds to Equation 2 is to follow the theory of (Anderson, 1982), use a neural network to approximate the score function and apply score matching. We notice that analytical form of diffusion bridge’s reverse transition kernel is directly obtainable, therefore the usage of maximum likelihood training can also be applied, which is performed by optimizing the Evidence Lower Bound (ELBO). It yields a more stable loss function compared with score matching and better aligns with the pretrained backbone. Given the specified terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
@@ 4.4.1 Maximum Likelihood Loss
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process and $\mathbb{D}_{KL}$ is the KL divergence. Utilizing Bayes’ theorem and Markov chain property, we can obtain:
@@
For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$. The mean value term is:
@@
$k_{1}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}$
$k_{2}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot m_{t-1}$
$k_{3}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
and the variance is:
@@
Assume $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ is also a Gaussian distribution parameterized as $\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ and let $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))=\tilde{\sigma}_{t}^{2}$ for derivation simplicity, maximizing the ELBO is equivalent to minimizing:
@@
Similar to the strategy in DDPM (Ho et al., 2020), we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as:
$\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right],$ (9)
4.4.2 Image Consistency Loss
In addition to the likelihood loss evaluated in latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
@@ 4.4.2 Image Consistency Loss
The reason is that the compression of input data in the latent space will lead to a loss of high-frequency information, harming the model’s ability of preserving fine-grained spatial details. By introducing an auxiliary loss at image level, we ensure that the model is capable of retaining spatial information which is essential for precise per-pixel prediction. This dual loss strategy leverages the compactness of latent space for efficiency while still anchoring the training process to the fidelity of full-resolution dense map prediction, thus leading to performance improvement. Combined these two losses together, our final training objective becomes:
@@
Algorithm 1 Training
1:  repeat
2:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
11:  until Converged
Algorithm 2 Inference
1:  $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
2:  for $t=T,\dots,1$ do
7:  end for
8:  $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
9:  return $\mathbf{y}$

### v3

@@ Abstract
Diffusion models demonstrate remarkable capabilities in capturing complex data distributions and have achieved compelling results in many generative tasks. While they have recently been extended to dense prediction tasks such as depth estimation and surface normal prediction, their full potential in this area remains under-explored. In dense prediction settings, target signal maps and input images are pixel-wise aligned. This makes conventional noise-to-data generation paradigm inefficient, as input images can serve as more informative prior compared to pure noise. Diffusion bridge models, which support data-to-data generation between two general data distributions, offer a promising alternative, but they typically fail to exploit the rich visual priors embedded in large pretrained foundation models. To address these limitations, we integrate diffusion bridge formulation with structured visual priors and introduce DPBridge, the first latent diffusion bridge framework for dense prediction tasks. Our method presents three key contributions: (1) a tractable reverse transition kernel for diffusion bridge process, enabling maximum likelihood training scheme for better compatibility with pretrained backbones; (2) a distribution-aligned normalization technique to mitigate the discrepancies between the bridge and standard diffusion processes; and (3) an auxiliary image consistency loss to preserve fine-grained details. Experiments across extensive benchmarks validate that our method consistently achieves superior performance, demonstrating its effectiveness and generalization capability under different scenarios.
@@ 1 Introduction
Dense prediction refers to a class of problems in which the goal is to assign a prediction value for every pixel in the input image. It encompasses a variety of fundamental tasks, including depth estimation [56, 39, 15], surface normal prediction [2, 22, 3], semantic segmentation [53, 8, 51], optical flow [49, 23, 54], etc.
Recently, image diffusion models have been applied to dense prediction tasks by framing them as conditional generation problems. Leveraging the ability to capture complex data distributions, diffusion models serve as powerful priors and yield promising results [28, 17, 16]. However, these approaches exhibit inherent limitations stemming from the underlying generative mechanism. Specifically, the denoising process always starts from Gaussian noise, which carries minimal information about the target signal distribution. In such cases, the distributional gap between initial Gaussian prior and target distribution remains large, necessitating more sampling steps and often reliance on manually designed post-hoc guidance [9, 46, 10] or customized generation pipeline [35]. This not only increases computational overhead but can also degrade performance and prolong convergence.
Instead of starting from noise, a more effective alternative is to directly initiate the generative process from the input. Since RGB images are pixel-wise aligned with their corresponding dense signal maps, they serve as informative priors. This motivates the exploration of diffusion bridge models [34, 37, 14, 33], another line of computational frameworks based on stochastic processes with fixed start and end states, designed to model probability flows between two given distributions. We argue that diffusion bridge models are more suitable for dense prediction tasks, as their data-to-data generation paradigm naturally leverage the structural correspondence between inputs and outputs, thereby mitigating inefficiencies inherent in conventional noise-to-data scheme. Despite theoretical feasibility, diffusion bridge-based methods [65, 31, 32] have not yet demonstrated competitive performance compared with standard diffusion counterparts. This performance gap largely stems from the fact that most bridge models are trained from scratch via score matching objectives, making them incompatible with commonly used pretrained diffusion backbone. As a result, they fail to exploit the embedded semantic and spatial prior knowledge, limiting their expressiveness and generalization ability in dense prediction settings.
Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing conditional probability transport between RGB images and corresponding dense maps in the latent space based on a tractable diffusion bridge process.
In this paper, we propose Dense Prediction Bridge (DPBridge), the first latent diffusion bridge model tailored for general dense prediction tasks by integrating the strengths of diffusion bridge modeling with visual priors from pretrained backbones. Specifically, we derive a fully tractable bridge process to connect the latent representations of RGB images and their corresponding dense signal maps, and formulate the generation process as a sequence of transformations that progressively synthesize the target maps from input data. Furthermore, we introduce finetuning strategies to better adapt pretrained diffusion backbones for DPBridge construction. Leveraging the rich visual priors not only enhances perceptual performance and generalization capability, but also significantly accelerates training convergence. Figure 1 provides a high-level overview of DPBridge framework and illustrates its effectiveness across three representative dense prediction tasks: depth estimation, surface normal prediction, and semantic segmentation. These results demonstrate that our framework is both theoretically general and practically adaptable to diverse dense prediction scenarios.
Contributions of this paper can be summarized as follows:
@@
  • We propose DPBridge, a generative framework to address dense prediction tasks by formulating them as instances of latent diffusion bridges.
  • We derive a tractable reverse transition kernel for diffusion bridge process, enabling maximum likelihood training scheme. This is compatible with pretrained backbones, thus improving both training stability and the exploitation of visual prior knowledge.
  • We introduce a distribution-aligned normalization technique that mitigates distributional discrepancies between the bridge and standard diffusion processes, allowing for more effective finetuning. In addition, by incorporating auxiliary image consistency loss, we can better preserve fine-grained spatial details often lost in latent representations, enhancing per-pixel prediction accuracy.
@@
  • Our method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection.
@@ 2 Related Works
Dense Prediction. Dense predictions can be characterized as conditional distribution modeling and common solutions are categorized into discriminative and generative ones. Discriminative methods directly map input data to structured outputs in a feed-forward manner. [39] introduces transformer architecture and fusion mechanism to better capture both local and global context feature. [56] explore the usage of large-scale synthetic and pseudo-labeled images to enhance model’s capacity and robustness. [2] incorporates per-pixel ray directions and models the relative rotations between neighboring pixels. Recently, diffusion-based generative methods have been applied in this field and showed remarkable performance. [28] finetunes pretrained diffusion backbone, leveraging the rich visual knowledge within to enhance depth estimation. [16] takes advantage of diffusion priors to jointly estimate depth and surface normals from single image, enhancing detail preservation by decoupling complex data distributions into distinct sub-distributions. [19] exploits flow matching to facilitate the distribution shift between images and depth maps. Our work follow conventional generative pipeline, but instead of starting from pure noise, we establish direct connections between input image and target signal map distributions, and model conditional probability transport in iterative denoising manner.
@@
Diffusion and Bridge Models. Diffusion models [20, 47] are a family of generative models that are able to capture complex data distributions through a sequence of forward and reverse processes. They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image and video synthesis [41, 21], controlled editing [62, 4], inverse problem solving [9, 46], 3D assets generation [55, 63], and human-centered applications [26, 25, 24].
@@
Diffusion bridge models have gained increasing attention in generative modeling in recent years [37, 34, 65, 33, 14]. They offer a novel perspective for translating between distributions and enable a more effective data-to-data generation paradigm. Recent works have explored different types of bridge processes and their applications. [32, 31, 60] develop Schrödinger bridge, Brownian bridge, and Ornstein-Uhlenbeck bridge frameworks, applying them to image restoration tasks such as inpainting, super-resolution. [11] proposes a modified inference procedure that enforces data consistency. [7] introduces bridge models to text-to-speech systems, enhancing synthesis quality and sampling efficiency.
To the best of our knowledge, no prior work has explored the use of bridge models for dense prediction tasks to facilitate distribution transport between images and dense signal maps, despite their structural closeness. Our work aims to address this research gap. Conceptually, DPBridge is most related to DDBM [65], but the two differ in several key aspects: DDBM is trained from scratch using a score matching objective, whereas DPBridge is finetuned from large pretrained diffusion backbones using a maximum likelihood objective. In addition, we introduce training strategies tailored for efficient finetuning, enabling better use of visual priors and significantly improving performance on dense prediction settings.
@@ 3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that is conditioned to end at a specified terminal point at timestep $T$. One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints. We can use Doob’s h-transform technique [36, 42] to modify a reference process $\mathbb{Q}$ defined in Equation 1 to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens, where $\mathbf{x}$ is a predefined value. The conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ thus follows the law of
@@
$m_{t}=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right),\quad n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}},\quad\sigma_{t}=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}}$
@@ 4.1 Problem Formulation
We formulate different dense prediction tasks as instances of image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ denotes the target dense signal map (e.g. depth, surface normal, segmentation map), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input RGB image. In the following sections, we’ll elaborate the details of DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
@@ 4.2 Dense Prediction Bridge
Given paired training data, a straightforward approach is to construct the diffusion bridge model from scratch, following the formulation in Section 3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more, it lacks the ability to harness rich visual prior embedded in the foundation models pretrained with large-scale data,which have been shown to significantly enhance both perceptual quality and generalization capability in dense prediction tasks. To address these limitations and fully exploit the visual prior, our DPBridge is build upon pretrained diffusion backbones through finetuning. We explore several strategies to handle the misalignment between bridge process and standard diffusion process, facilitating more efficient knowledge transfer. An overview of our training pipeline is provided in Figure 2, followed by a detailed explanation of each component.
@@ 4.3 Latent Space Implementation
We follow common practice to perform diffusion steps in a compressed yet perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the dense signal map and RGB image are encoded using the VAE encoder, yielding latent representations $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, which serve as the initial and terminal points of the diffusion bridge process. In this way the diffusion bridge process is operated entirely in the latent space. To recover the original signal map, we use the VAE decoder to decode the predicted latent $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. For clarity, all subsequent derivations in this paper are conducted in latent space.
Figure 2: Overview of our training pipeline. We derive the maximum likelihood training objective of diffusion bridge process, propose a distribution-aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
@@ 4.4 Training Objective
Maximum Likelihood Loss. A widely adopted approach for learning the reverse process associated with Equation 2 follows the theoretical framework of [1], where a neural network is trained to approximate the score function [65]. In contrast, we observe that in the diffusion bridge setting, the reverse transition kernel admits a tractable analytical form. This insight enables us to adopt a maximum likelihood training paradigm in place of score matching. Since widely used latent diffusion models like Stable Diffusion [41] are pretrained under a likelihood-based loss function, finetuning with a consistent objective is more compatible, which helps preserve prior knowledge and improves training stability.
Specifically, the training process is performed by optimizing the Evidence Lower Bound (ELBO). Given the encoded training pair $\big{(}\mathbf{z}_{0},\mathbf{z}_{T}\big{)}$, the expectation of data log-likelihood possesses the following ELBO formulation:
@@
$\sum_{t=2}^{T}\mathbb{E}_{p(\mathbf{z}_{t}|\mathbf{z}_{0})}$ $\big{[}\mathbb{D}_{KL}\big{(}q(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\big{)}\big{]}\Big{]}$
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process. $\mathbb{D}_{KL}$ is the KL divergence. Utilizing Bayes’ theorem and Markov chain property, we can obtain:
@@
For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$. Based on Equation 3, we further derive the mean value term:
@@
$k_{1}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t},\quad k_{2}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot$ $m_{t-1},\quad k_{3}=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
and the variance term:
@@
Assuming $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})\sim\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ is also a Gaussian distribution with fixed variance $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)=\tilde{\sigma}_{t}^{2}$ for simplicity of derivation. Following the strategy of DDPM [20], we reparameterize the model $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to predict noise instead of posterior mean. Under this formulation, the ELBO objective is equivalent to:
@@
$\cong\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
Image Consistency Loss. In addition to the latent-space likelihood loss, we also incorporate an image-level consistency loss, where we first decode the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE.
@@
This additional supervision addresses the loss of high-frequency structures caused by latent-space compression, which can impair the model’s ability to preserve fine-grained details. By introducing this auxiliary image-space loss, we encourage the model to retain spatial information critical for accurate per-pixel prediction. This dual-loss strategy combines the efficiency of latent-space modeling with the fidelity of image-space supervision, leading to improved performance. The final training objective is given by: